an ideal gas at 50 psig and 110 f is heated in a closed cylindrical vessel. what is temperature fi the pressure gauge reading is 100 psig


1
Expert's answer
2021-03-05T02:52:37-0500

According to the combined gas law:

p1/t1 = p2/t2

where p1 - initial pressure, p2 - final pressure, t1 - initial temperature, t2 - final pressure

From here:

t2 = p2t1 / p1 = 100 psig × 110 f / 50 psig = 220 f


Answer: 220 f